OpenStudy (anonymous):

i need help solving x2 -4x +3 =0 by completing the square

OpenStudy (anonymous):

x^2-4x=-3 x^2-4x+4=-3+4 (x-2)^2=1 x-2=+/- 1

OpenStudy (anonymous):

\[x ^{2} - 4x + 3 = 0\] x^2 - 4x = -3 x^2 -4x + 4 = -3 +4 (x-2)^2 = 1 x - 2 = + or - 1

OpenStudy (anonymous):

thanks